OpenStudy (eric_d):

Express it in partial fraction 15) 1/x^2(x^-1) = @ganeshie8

ganeshie8 (ganeshie8):

check the function again, looks there is some typo..

OpenStudy (eric_d):

1/x^2(x^-1)

OpenStudy (eric_d):

= A/x + B/x^2 + C/(x+1) + D/(x-1) Do you know this method..? Can u teach me how to make all the denominators same?

ganeshie8 (ganeshie8):

that `^` is throwing me off :/

OpenStudy (eric_d):

sorry

ganeshie8 (ganeshie8):

your function is like below :\[\large \dfrac{1}{x^2(x-1)}\] ?

OpenStudy (eric_d):

jst realised it! it's 1/x^2(x^2-1)

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{x^2(x^2-1)}\]

OpenStudy (eric_d):

yes..

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{x^2(x^2-1)} = \dfrac{1}{x^2(x+1)(x-1)}\]

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{x^2(x+1)(x-1)} = \dfrac{A}{x} + \dfrac{B}{x^2} + \dfrac{C}{x+1}+ \dfrac{D}{x-1}\]

ganeshie8 (ganeshie8):

so your setup looks good so far :)

ganeshie8 (ganeshie8):

your goal is to find those constants A,B,C,D

OpenStudy (eric_d):

I want to clarify something...

OpenStudy (eric_d):

If

ganeshie8 (ganeshie8):

sure, ask..

OpenStudy (eric_d):

Thnks. If there's another question like 1/x^3(x+1)(x-1)...How wld the setup be

ganeshie8 (ganeshie8):

you can guess i hope :)

ganeshie8 (ganeshie8):

\[\dfrac{1}{x^3(x+1)(x-1)} = \dfrac{A}{x} + \dfrac{B}{x^2} +\dfrac{C}{x^3} + \dfrac{D}{x+1}+ \dfrac{E}{x-1} \]

ganeshie8 (ganeshie8):

adds up one more constant ^

OpenStudy (eric_d):

Okay.

OpenStudy (eric_d):

Let's continue wth the first question..

ganeshie8 (ganeshie8):

yeah familiar with `Cover up` method ?

OpenStudy (eric_d):

I'm nt sure. You can show it to me..

ganeshie8 (ganeshie8):

its okay, lets work it using the method known to you : comparing coefficients

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{x^2(x+1)(x-1)} = \dfrac{A}{x} + \dfrac{B}{x^2} + \dfrac{C}{x+1}+ \dfrac{D}{x-1}\]

ganeshie8 (ganeshie8):

multiply the right hand side by a special kind of \(\large \color{Red}{1}\) : \(\large \color{red}{ \dfrac{x^2(x+1)(x-1)}{x^2(x+1)(x-1)}}\)

OpenStudy (eric_d):

after that

ganeshie8 (ganeshie8):

\[ \dfrac{1}{x^2(x+1)(x-1)} = \large \color{red}{ \dfrac{x^2(x+1)(x-1)}{x^2(x+1)(x-1)}} \left[\dfrac{A}{x} + \dfrac{B}{x^2} + \dfrac{C}{x+1}+ \dfrac{D}{x-1}\right]\]

ganeshie8 (ganeshie8):

\[ \dfrac{1}{x^2(x+1)(x-1)} = \large \color{red}{ \dfrac{1}{x^2(x+1)(x-1)}}\\ \left[A\color{red}{x(x+1)(x-1)} + \\ B\color{red}{(x+1)(x-1)} + \\ C\color{red}{x^2(x-1)} \\D\color{red}{x^2(x+1)} \right]\]

ganeshie8 (ganeshie8):

canceling denominator both sides gives : \[1 = Ax(x+1)(x-1) +B(x+1)(x-1) + Cx^2(x-1) + Dx^2(x+1)\]

ganeshie8 (ganeshie8):

see if that looks okay so far

OpenStudy (eric_d):

How to get this... Ax(x+1)(x-1)

ganeshie8 (ganeshie8):

after distributing, the first termis : \(\large \color{Red}{x^2(x+1)(x-1)}\dfrac{A}{x}\)

ganeshie8 (ganeshie8):

right ?

OpenStudy (eric_d):

yes ... So, I need to distribute each constant

ganeshie8 (ganeshie8):

yes, we do that to get rid off of the fractions

ganeshie8 (ganeshie8):

notice that the denominator x cancels out

ganeshie8 (ganeshie8):

\(\large \color{Red}{x^{\not 2}(x+1)(x-1)}\dfrac{A}{\not x}\)

ganeshie8 (ganeshie8):

leaving you with \( \large A \color{Red}{x(x+1)(x-1)} \)

OpenStudy (eric_d):

what's nxt

ganeshie8 (ganeshie8):

compare the coefficients both sides

ganeshie8 (ganeshie8):

\[1 = Ax(x+1)(x-1) +B(x+1)(x-1) + Cx^2(x-1) + Dx^2(x+1)\]

ganeshie8 (ganeshie8):

or may be plugin some values

ganeshie8 (ganeshie8):

plugin x = 1, and see what you get

OpenStudy (eric_d):

Constant A,B,C = 0

OpenStudy (eric_d):

1=2D D=1/2

OpenStudy (eric_d):

What do I need to do after getting the values of all the constant

ganeshie8 (ganeshie8):

Yes ! next think of another good value to plugin

OpenStudy (eric_d):

okay

ganeshie8 (ganeshie8):

after getting all the constants, you're done !

ganeshie8 (ganeshie8):

The goal of partial fractions is to split the given rational functions into separate fractions, thats all.

OpenStudy (eric_d):

Understood!

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{x^2(x+1)(x-1)} = \dfrac{A}{x} + \dfrac{B}{x^2} + \dfrac{C}{x+1}+ \dfrac{D}{x-1}\]

OpenStudy (eric_d):

Thnx @ganeshie8

ganeshie8 (ganeshie8):

Once you have the values of all the constants, just plug them above ^ and conclude !

OpenStudy (eric_d):

when x=1 D=1/2 when x=-1 C=-1/2 How to find A and B

OpenStudy (anonymous):

ganshie is currently offline that your problem

OpenStudy (anonymous):

but just use the formula he gave u to plug in the numbers

OpenStudy (anonymous):

im sorry but i am no that good at math but i am your go to person for history

OpenStudy (eric_d):

By the way, the topic is about partial fraction

OpenStudy (eric_d):

Okay, I'll wait for Ganesh coz I've nt learn matrix method to solve such questiions. But, u can show ur working here using matrices